Write a quadratic equation having the given numbers as solutions -7 and -2
Found 2 solutions by Fombitz, checkley77:
Answer by Fombitz(32388)
Answer by checkley77(12844)

X^2+9X+14=0
(X+7)(X+2)=0
X=7=0
X=-7 ANS.
X+2=0
X=-2 ANS.